OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 3x+4y=17 -4x-3y=-18 please help

OpenStudy (shamim):

i think ur second equation is incorrect

OpenStudy (anonymous):

you first want to mult the first by 4 and the second by 3, this will allow you to cancel out the x values when u add and solve for y

OpenStudy (anonymous):

ohh my bad the second equation should be -4x-3y=-18

OpenStudy (shamim):

sevond equation has no x

OpenStudy (anonymous):

i fixed it

OpenStudy (anonymous):

could either of you go further into the question im still not getting it

OpenStudy (anonymous):

did you mult the two equations already?

OpenStudy (anonymous):

i am right now and its just not clicking so far

OpenStudy (anonymous):

4(3x+4y=17) = 12x + 16y = 68 3(-4x-3y=-18) = -12x -9y = -54

OpenStudy (anonymous):

when u add, the x values cancel each other out

OpenStudy (anonymous):

7y=14

OpenStudy (anonymous):

divide both sides by 7 to get y alone and u will see that y equals 2

OpenStudy (anonymous):

then u plug 2 into one of the equations and solve the equation for x

OpenStudy (anonymous):

the ordered pair goes (x, y)

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

kinda but im not getting to an answer

OpenStudy (anonymous):

for which part

OpenStudy (anonymous):

what are you getting and the final answer

OpenStudy (anonymous):

well, after we find that y=2 we plug that in

OpenStudy (anonymous):

3x +4(2)=17

OpenStudy (anonymous):

3x+8=17

OpenStudy (anonymous):

3x=9

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

ordered pair (3,2)

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes thanks